# Merlin: A Vision Language Foundation Model for 3D Computed Tomography

### v1

@@ 4.2 Vision-Language-EHR Model
Model Training: We use binary cross entropy for the phenotype classification loss and InfoNCE Oord et al. (2018); Radford et al. (2021a) loss for contrastive learning with the radiology reports. We use an AdamW Loshchilov and Hutter (2017) optimizer with an initial learning rate of 1e-5, betas=$(0.9,0.999)$, and a cosine learning rate scheduler with number of epochs for decay to 0 set to 300. We use gradient checkpointing for both the visual and text encoders and train with FP16 mixed precision. This allows us to maximize a batch size of 18 on a single 48GB A6000 GPU.
@@
Data Splits: We divide the pretraining dataset into splits of size 60% (15,331 CTs) for training, 20% (5,060 CTs) for validation, and 20% (5,137 CTs) for testing. We ensure that multiple CTs from a single patient exist in a single split. The external dataset, from another university medical center, consists of 7,000 CTs which are used for testing.
@@ 4.3 Evaluations
We select evaluation tasks consisting of non-adapted tasks, which Merlin can perform out of the box without additional adaptation. These tasks include zero-shot findings classification, phenotype classification, and zero-shot cross-modal retrieval. We also select tasks that require adapting Merlin, which include 5-year disease prediction, radiology report generation, and 3D semantic segmentation. We select both non-adapted tasks and adapted tasks to demonstrate Merlin’s effectiveness without fine-tuning, as well as its adaptability when fine-tuned for specific applications.
@@ 4.4 Non-Adapted Tasks
Phenotype Classification: To generate the labels for this task, we group ICD-9 and ICD-10 codes into phenotypes using the PheWAS Denny et al. (2013) phecode mapping. We compare Swin Transformer, ConvNeXt, and ResNet backbone architectures (Figure 3(c)), as well as various model stem hyper-parameters (Figure 3(d)). See section 4.2 for additional details. We report average AUROC and AUPRC using all phenotypes that have more than 20 positive examples in the test set, in order to ensure a meaningful measure of performance. We additionally assess how phenotype classification performance varies across the pretraining dataset sizes of 1%, 10%, 20%, 40%, and 100% of 15,331 total pretraining samples.
Counterfactual analysis (Figure 3(e)) seeks to evaluate whether a model is using expected features during phenotype classification. Models can instead learn shortcuts from spurious correlations in the training data Ross et al. (2017); Geirhos et al. (2020); Cohen et al. (2021b), which are easier to learn compared to more subtle physiological features that the model should use. We employ the Latent Shift approach Cohen et al. (2021a) for counterfactual generation. This method uses a latent variable model, which represents the input image in a low dimensional space and reconstructs the image in its output. The gradient of the Merlin prediction is passed back to the low-dimensional space, where the latent representation is modified such that Merlin’s prediction decreases. This generates a modified image at the output of the latent variable model, which can be observed. The latent variable model regularizes the change in pixels so that the image remains close to the data manifold. The result of the Merlin counterfactual generation process is a new CT volume, which is evaluated by observing how the classifier prediction changes and how the image features driving the classification output are exaggerated or curtailed. We leverage this method to qualitatively verify that the features used during classification are consistent with what we expect the model to have learned. For this evaluation, we identify pleural effusion and splenomegaly as phenotypes with features that are well understood and can be observed in 2D slices. We select the presented volumes such that the phenotype prediction is above the decision boundary determined by AUROC curve analysis, to ensure that we are explaining a positive prediction.
@@ 4.5 Adapted Tasks
3D Semantic Segmentation To adapt Merlin for segmentation, we add a UNet Ronneberger et al. (2015) decoder and skip connections between the Merlin encoder and the decoder. Each block of the decoder consists of a 3D transpose convolution operation, with a kernel size of 2 and a stride of 2 in each dimension, followed by two 3D convolutions, with kernel sizes of 3 and strides of 1 in each dimension. Each of the two 3D convolution operations is followed by a 3D batchnorm and ReLU activation. We also add skip connections where outputs from the Merlin ResNet blocks are concatenated with outputs from the 3D transpose convolution in the decoder before being passed to the subsequent 3D convolutional layers. For model training, we use both a cross-entropy loss and a Dice loss. We use an AdamW optimizer with betas=$(0.9,0.999)$ and a learning rate of 5e-5. To mirror pretraining, we segment full volumes of size 224x224x160. We achieve this with 4 gradient accumulation steps and a local batchsize of 1, providing a total batchsize of 4.
@@
We compare performance across several model variations using Dice score. We compare Merlin to models with the same architecture but initialized randomly or using ImageNet I3D initialization. We also compare against, Swin UNETR Tang et al. (2022), an architecture specifically designed for 3D medical image segmentation.
@@ Acknowledgements
AC receives research support from the National Institutes of Health (grants - R01 HL167974, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060, and contracts 75N92020C00008, 75N92020C00021); and from GE Healthcare, Philips, Amazon, Microsoft/OpenAI, and Stability.ai.
@@
Contributions L.B. collected data, developed code, trained models, ran experiments, analyzed results, created figures, and wrote the manuscript. All authors reviewed the manuscript and provided meaningful revisions and feedback. J.P.C, A.K., D.V.V, M.P., Z.C., J.B.D, E.R., C.B., M.E.K.J, S.O., M.V., J.M.J.V., Z.N., D.A., W.W., S.G., and A.S.C. provided technical advice. J.P.C developed the counterfactual generation method for CT and ran counterfactual experiments. A.K. carried out model evaluations on external datasets. D.V.V. assisted in collecting zero-shot evaluation labels and facilitated radiologist annotations of generated reports. S.J.S.G. ran model inference on the external clinical dataset. M.P., Z.C., J.B.D., E.R., C.T., E.A.J. assisted with model evaluations. Z.H. and J.F. assisted with dataset anonymization. C.B., E.A.J, A.J., R.D.B., A.W., C.P.L., J.H., and S.G., provided clinical input and feedback. C.B. provided counterfactual annotations. S.G. and C.B. provided annotations of generated reports. A.S.C. guided the project, serving as principal investigator and advising on technical details and overall direction. No funders or third parties were involved in study design, analysis or writing.
Supplementary
@@
$Encoder$ $Init$ $Labels$ $Stem$ All Phecodes Upper Quartile Lower Quartile
@@
$Init$ $Labels$ $Split$ $Stem$ All Phecodes Upper Quartile Lower Quartile
$Text$ KSz/ N=691, Prev=3.1% by Prevalence by Prevalence
@@
$Method$ Average F1 Score
OpenCLIP (Internal) .276 [.262, .288]
BioMedCLIP (Internal) .285 [.274, .295]
Merlin (Internal) .741 [.727, .755]
Merlin (External) .647 [.607, .678]
Merlin (VerSe) .767 [.630, .867]
$Init$ $Labels$ $Split$ Average F1 Score
$Text$
I3D Report .730 [.714, .744]
I3D Staged .669 [.653, .683]
I3D Staged .735 [.719, .748]
I3D MTL .656 [.640, .671]
Rand MTL .698 [.681, .711]
I3D MTL .741 [.727, .755]
$Task$ $Method$ Recall@1 Recall@8

### v2

@@ 4.2 Vision-Language-EHR Model
Model Training: We use binary cross entropy for the phenotype classification loss and InfoNCE Oord et al. (2018); Radford et al. (2021a) loss for contrastive learning with the radiology reports, where the embeddings for both the image and text encoders in our implementation of the InfoNCE loss had a dimension of 512. This matches the embedding dimension used in the ViT-Base model in OpenCLIP experiments Cherti et al. (2023). We use an AdamW Loshchilov and Hutter (2017) optimizer with an initial learning rate of 1e-5, betas=$(0.9,0.999)$, and a cosine learning rate scheduler with number of epochs for decay to 0 set to 300. We use gradient checkpointing for both the visual and text encoders and train with FP16 mixed precision. This allows us to maximize a batch size of 18 on a single 48GB A6000 GPU.
@@
Data Splits: We divide the pretraining dataset into splits of size 60% (15,331 CTs) for training, 20% (5,060 CTs) for validation, and 20% (5,137 CTs) for testing. We ensure that multiple CTs from a single patient do not exist in a single split. The external dataset, from another university medical center, consists of 7,000 CTs which are used for testing. The external dataset from out-of-state site 1 consisted of 6,997 abdominal CT scans used for testing. In-state external sites 2 and 3 included 25,997 and 4,872 abdominal CT scans, respectively, used for testing; site 3 additionally included 6,243 chest CT scans used for testing.
@@ 4.3 Evaluations
We select evaluation tasks consisting of non-adapted tasks, which Merlin can perform out of the box without additional adaptation. These tasks include zero-shot findings classification, phenotype classification, and zero-shot cross-modal retrieval. We also select tasks that require adapting Merlin, which include 5-year disease prediction, radiology report generation, and 3D semantic segmentation. We select both non-adapted tasks and adapted tasks to demonstrate Merlin’s effectiveness without fine-tuning, as well as its adaptability when fine-tuned for specific applications. To comprehensively evaluate Merlin’s performance against alternative architectures, we compare Merlin to finetuned 2D VLMs, 2D-to-3D lifted VLMs, and 3D vision-only models in data-scarce and fully-supervised settings.
@@ 4.4 Non-Adapted Tasks
Phenotype Classification: To generate the labels for this task, we group ICD-9 and ICD-10 codes into phenotypes using the PheWAS Denny et al. (2013) phecode mapping. We compare Swin Transformer, ConvNeXt, and ResNet backbone architectures (Figure 3c), as well as various model stem hyper-parameters (Figure 3d). See section 4.2 for additional details. We report average AUROC and AUPRC using all phenotypes that have more than 20 positive examples in the test set, in order to ensure a meaningful measure of performance. We additionally assess how phenotype classification performance varies across the pretraining dataset sizes of 1%, 10%, 20%, 40%, and 100% of 15,331 total pretraining samples.
@@ 4.5 Adapted Tasks
3D Semantic Segmentation: To adapt Merlin for segmentation, we add a UNet Ronneberger et al. (2015) decoder and skip connections between the Merlin encoder and the decoder. Each block of the decoder consists of a 3D transpose convolution operation, with a kernel size of 2 and a stride of 2 in each dimension, followed by two 3D convolutions, with kernel sizes of 3 and strides of 1 in each dimension. Each of the two 3D convolution operations is followed by a 3D batchnorm and ReLU activation. We also add skip connections where outputs from the Merlin ResNet blocks are concatenated with outputs from the 3D transpose convolution in the decoder before being passed to the subsequent 3D convolutional layers. To mirror pretraining, we segment full volumes of size 224x224x160.
@@
We evaluated various network architectures and initialization strategies within the nnUNet framework. Specifically, we compared the nnUNet 3D full-resolution architecture with other configurations, including a Swin UNETR Tang et al. (2022), an architecture specifically designed for 3D medical image segmentation, and a Resnet152 encoder-decoder architecture initialized with random, Imagenet, and Merlin weights (Figure 7a). The training setup followed the default nnUNet configuration, including data augmentation steps, a 1000-epoch training schedule, and a polynomial learning rate scheduler. We evaluated the model in both 10% and 100% training data domains using Dice score.
4.6 Alternative Architecture Experiments
2D VLMs: We trained three 2D vision-language foundation models: a 2D Resnet152 vision encoder paired with a clinical longformer text encoder, OpenCLIP Cherti et al. (2023) finetuned on abdominal CT data, and BiomedCLIP Zhang et al. (2023b) finetuned on abdominal CT data (Figure 8a). Each model used MTL, incorporating EHR phenotypes and radiology reports split into anatomical sections, to ensure a comparable training approach to the 3D Merlin model. This setup finetuned both the image and text encoders using the EHR and a contrastive loss head. During training, we randomly selected a single slice from each 3D volume as input. The Resnet152 model was initialized with ImageNet weights. All models were finetuned with a batch size of 128, 100 epochs, using 100% of the training data.
2D-to-3D lifted VLMs: To adapt the aforementioned 2D vision-language models for 3D data, we finetuned the models by processing each slice of the abdominal CT volume individually and concatenating them using a learned projection layer (Figure 8b). Instead of naive averaging, we propose transforming the features from each highly-redundant 2D slicewise embeddings from the 3D volume into an unified embedding through this learned layer. More specifically, the 160 slice embeddings are processed through a sequential linear transformation, first reducing them to 64 embeddings with a ReLU non-linearity, and then further compressed into a single embedding of size 512. The image encoder and classification heads from the VLM were then finetuned for 10 epochs on 100% of the training data, with maximal attainable batch sizes of 2, 8, and 16 for 2D Resnet152, OpenCLIP, and BiomedCLIP, respectively.
3D vision-only models: We trained two state-of-the-art 3D vision-only foundation models: Swin UNETRHatamizadeh et al. (2022) and Inflated 3D (I3D) Resnet152, using a self-supervised learning (SSL) pretraining strategy Tang et al. (2022) (Figure 8c). The SSL pretraining process involved randomly cropping abdominal CT volumes into subvolumes, followed by augmentations such as random cutout and rotation. These augmented subvolumes were then fed into the models, with pretraining tasks including inpainting, contrastive learning, and rotation prediction to learn meaningful image representations, based on prior work Tang et al. (2022). For Swin UNETR, pretraining was performed with a batch size of 1, a sliding window size of 2, and 50K training steps. For I3D Resnet152, the model utilized convolutional embeddings with convolutional transpose layers for the L1 reconstruction loss. Finetuning the I3D Resnet152 was conducted with a batch size of 3, a sliding window size of 2, and 50K training steps.
4.7 Alternative Architecture Evaluation
To ensure a fair comparison between the vision-only and vision-language baselines, we evaluate the models on two tasks: findings-based disease classification and EHR phenotype classification. We here describe the evaluation tasks, which involve training linear-probes while keeping the image encoder frozen. This setup allows us to compare pretrained image representations in isolation and compare Merlin to recent CT embedding foundation modelsCodella et al. (2024); Yang et al. (2024). Full finetuning results, where both the image encoder and classification heads are trained, are presented in Supplemental Section C.1. For the 3D vision-only models, the embedding dimensions were 768 for SwinUNETR and 2048 for I3D Resnet152. For the 2D and the 2D-to-3D VLMs, the embedding dimensions were 768 for BiomedCLIP, 2048 for ResnetCLIP, and 1024 for OpenCLIP. For the CT embedding foundation models, MedImageInsight used an embedding dimension of 1024, while Google CT Foundation used an embedding dimension of 1408. Details on processing the Merlin dataset through these CT embedding foundation models are provided in Supplemental Section C.1.
Findings-based disease classification task: The findings-based disease classification task was adapted from the zero-shot findings classification task (Section 4.4), where the 30 findings originally used for zero-shot evaluation were reformulated for supervised learning. We parsed the radiology reports and categorized each finding as positive, negative, or missing from the report. While most conventional strategies involve discarding missing labels during training, we adopted a more inclusive strategy to use all our data. We framed the task as a multilabel classification task, introducing thirty separate binary heads, one for each finding. Image embeddings were used as input to the multihead linear classifiers, and the loss was masked for subjects missing a specific finding before backpropagationXue et al. (2024); Blankemeier et al. (2022). Linear probes were trained from scratch using Kaiming initialization with the AdamW optimizer, an initial learning rate of 1e-5, and a cosine decay scheduler (300 epochs to zero learning rate), for a total of 10 epochs using binary cross-entropy loss on the abdominal CT dataset (15,331 scans for training and 5,060 for validation). We evaluated the performance of these baselines using 10% and 100% of the training data. This approach standardizes image encoder evaluation across various model architectures, including 3D vision-only foundation models, 2D vision-language models, 2D finetuned vision-language models, 2D-to-3D lifted vision-language models, MedImageInsight, Google CT foundation model, and Merlin.
Phenotype classification For the EHR phenotype classification task, we train a linear probe that maps the embeddings to 1,692 target labels. The probe consists of a single linear layer with input dimensionality matching the model’s image encoder embedding dimensions and an output of 1,692 classes. We use the same optimizer, scheduler, training hyperparameters, and image dataset as in the multihead linear probe experiments (findings-based disease classification task). All linear probes were trained from scratch using Kaiming initialization to maintain consistent comparisons across models. Similar to the findings-based disease classification task, we finetuned the models on 10% and 100% of the training data to assess classification performance across different dataset sizes.
4.8 External Validation Experiments
We evaluate Merlin on 44,098 external CT scans, comprising 37,855 abdominal CTs from three external sites [Site #1: n = 6,997, out-of-state site; Site #2: n = 25,986, in-state site; Site #3: n = 4,872, in-state site] and 6,243 chest CTs from Site #3.
Zero-shot findings classification: We evaluate Merlin on the zero-shot findings classification task and compare Merlin against alternative architecture baselines: 2D/2D-to-3D BiomedCLIP, OpenCLIP, and ResnetCLIP. All models were fine-tuned on the Merlin dataset in a multi-task manner, analogous to Merlin’s training. This setup ensured that observed differences reflected architectural paradigms (2D versus 2D-to-3D versus full 3D [Merlin]) rather than differences in pretraining data quality. Ground-truth labels for zero-shot classification were derived by matching positive and negative prompts to the findings sections of radiology reports using regular expressions. The same positive and negative prompts as in the internal test set (Section  4.4) were used. For the External Site #3, labels were generated using the Qwen3-30B LLM Yang et al. (2025) applied to the findings section of their abdominal CT reports.
Chest CT linear probe evaluation: We leveraged a dataset of 63,711 chest CTs from External Site #3, comprising 57,468 scans for training and 6,243 for testing. We specifically only trained a linear probe on the frozen Merlin image encoder to assess the quality of the baseline Merlin embeddings. Each scan is paired with a radiology report, and labels were derived from the findings section using the Qwen-30B LLM Yang et al. (2025) to identify 24 conditions: lung nodules, consolidation, ground-glass opacities, atelectasis, interstitial thickening, emphysema, bronchiectasis, pulmonary embolism, pulmonary edema, lung cavitations, pleural effusion, pneumothorax, pleural thickening, lymphadenopathy, dilated aorta, aortic aneurysm, aortic dissection, atherosclerosis, pericardial effusion, metastases, fracture, hepatic cysts, liver mass, and hiatal hernia. We trained only a linear probe using balanced logistic regression to correct for skewed disease prevalence, without explicit resampling. The implementation was based on the scikit-learn library, mapping each model’s pretrained embeddings to the extracted labels. From a model comparison perspective, we evaluated Merlin against recent chest CT foundation models—CT-CLIP Hamamci et al. (2024b), M3FM Niu et al. (2025), and CT-FM Pai et al. (2025)—by training a separate linear probe for each disease and model.
@@ Acknowledgements
A.S.C. receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB 027060 and P50 HD118632; the Advanced Research Projects Agency for Health (ARPA-H) Biomedical Data Fabric (BDF) and the Chatbot Accuracy and Reliability Evaluation (CARE) programmes (contracts AY2AX000045 and 1AYSAX0000024-01); and the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the ARPA-H. C.B. receives research support from the Promedica Foundation.
@@
L.B. and A.K. collected data, developed code, trained models, ran experiments, analysed results, created figures and wrote the manuscript. All authors reviewed the manuscript and provided revisions and feedback. J.P.C., A.K., D.V.V., M.P., Z.C., J.-B.D., E.R., R.H., C.B., M.E.K.J., S.O., M.V., J.M.J.V., Z.F., Z.N., D.A., W.-H.W., S.G. and A.S.C. provided technical advice. J.P.C. developed the counterfactual generation method for CT scans and ran counterfactual experiments. A.K. carried out model evaluations on external datasets. J.L. developed a Merlin dataset inference pipeline for the Google CT. D.V.V. assisted in collecting zero-shot evaluation labels and facilitated radiologist annotations of generated reports. S.J.S.G., H.Y. and A.W. ran model inference on external clinical dataset 1. L. Liu, L. Lian, Y.W. and A.Y. ran model inference on external clinical dataset 3. M.P., Z.C., J.-B.D., E.R., C.T. and E.A.J. assisted with model evaluations. Z.H. and J.F. assisted with dataset anonymization. C.B., E.A.J., N.A., G.Z., M.W., A.J., R.D.B., A.W., C.P.L., M.W., J.H. and S.G. provided clinical input and feedback. C.B. provided counterfactual annotations. S.G. and C.B. provided annotations of generated reports. N.H.S., C.P.L., S.G. and A.S.C. provided research support for the project. A.S.C. guided the project, serving as principal investigator and advising on technical details and overall direction. No funders or third parties were involved in study design, analysis or writing.
Appendix A Extended Methods
A.1 Dataset Details
@@
Appendix B Extended Results
B.1 Zero-shot Classification
MethodMethod Average F1 Score
OpenCLIP (Internal) .276 [.262, .288]
BioMedCLIP (Internal) .285 [.274, .295]
Merlin (Internal) .741 [.727, .755]
Merlin (External) .647 [.607, .678]
Merlin (VerSe) .767 [.630, .867]
InitInit LabelsLabels SplitSplit Average F1 Score
TextText
I3D Report .730 [.714, .744]
I3D Staged .669 [.653, .683]
I3D Staged .735 [.719, .748]
I3D MTL .656 [.640, .671]
Rand MTL .698 [.681, .711]
I3D MTL .741 [.727, .755]
Scaling Law Experiments: We expand Figure 2d by providing a quantitative analysis of how increasing the pretraining dataset affects zero-shot performance. This analysis allows us to estimate the amount of pretraining data needed to achieve a given zero-shot classification performance. The resulting power law is $\overline{F1}=0.458D^{0.0524}$, where $D$ represents the number of paired pretraining CT imaging and report data.
Additional 2D and 2D-to-3D Experiments: We expanded our analysis to present the top-k pooling results for existing baseline models and included the zero-shot findings evaluation to encompass finetuned 2D and 2D-to-3D lifted VLMs (Figure 10). We observed that top-k pooling does increase F1 score compared to mean-pooling.
We first evaluated off-the-shelf 2D BiomedCLIP and OpenCLIP models, without finetuning, by calculating average F1 scores for top-k pooling (k=1, 10, 50). Our findings indicate that top-k pooling improves F1 scores for zero-shot evaluation compared to average pooling, with performance decreasing as k increases (Figure 10a). For instance, when comparing k=1 to the average (AVG), BiomedCLIP shows a 32.0% increase in F1 score, while OpenCLIP shows a 29.0% increase in F1 score.
We repeated the top-k pooling experiments after 2D finetuning of BiomedCLIP, OpenCLIP, and ResnetCLIP (Figure 10b). Interestingly, after finetuning, pooling across averaged slices yielded the highest F1 scores. When comparing average pooling F1 scores between the finetuned and non-finetuned versions of BiomedCLIP and OpenCLIP, we observed that finetuning increased BiomedCLIP’s F1 score by 52.0%, whereas OpenCLIP’s F1 score decreased by 2.1%. For instance, with the finetuned models, average pooling outperformed K=1 pooling, with BiomedCLIP showing a 4.9% increase in F1 score, OpenCLIP a 5.0% increase, and ResnetCLIP a 22.0% increase. When comparing the finetuned 2D baselines on average pooling, BiomedCLIP achieved the highest F1 score of 0.432 (95% CI [0.419-0.443]), while OpenCLIP and ResnetCLIP demonstrated lower F1 scores of 0.272 (95% CI [0.260-0.283]) and 0.293 (95% CI [0.279-0.305]), respectively. This trend may be attributed to the 2D finetuning process, which involved selecting a random slice for training. By minimizing reliance on specific slices, this approach reduces potential advantages from pooling. To assess potential overfitting during 2D finetuning, we examined the model’s performance across training epochs by tracking changes in zero-shot F1 score. The analysis revealed minimal signs of overfitting, with training performance saturating around 30 epochs (Figure 10d).
Finally, we compared the 2D-to-3D lifted VLMs to Merlin on the zero-shot findings classification task. Among the aggregation strategies, 2D-to-3D OpenCLIP achieved the highest F1 score of 0.632 (95% CI [0.613-0.647]), followed by 2D-to-3D ResnetCLIP with an F1 score of 0.614 (95% CI [0.595-0.623]) and BiomeCLIP with 0.587 (95% CI [0.568-0.604]) (Figure 10c). Despite these results, OpenCLIP’s F1 score was still 17.0% lower than that of Merlin.
Figure 10: Zero-shot findings classification extended. (a) Top-k pooling on existing 2D BiomedCLIP and OpenCLIP baselines. (b) Top-k pooling on 2D finetuned BiomedCLIP, OpenCLIP, and ResnetCLIP baselines. Finetuning involved training both the image and text encoders following the strategy outlined in Figure 8b. (c) 2D-to-3D lifted VLM zero-shot evaluation. Merlin tends to perform better on this task compared to baselines from other training paradigms. (d) BiomedCLIP’s F1 on zero-shot findings classification task across training epochs.
B.2 Phenotype Classification
EncoderEncoder InitInit LabelsLabels StemStem All Phecodes Upper Quartile Lower Quartile
@@ B.2 Phenotype Classification
InitInit LabelsLabels SplitSplit StemStem All Phecodes Upper Quartile Lower Quartile
TextText KSz/ N=691, Prev=3.1% by Prevalence by Prevalence
@@
Scaling Law Experiments: We expand Figure 3b by computing data scaling law curves and performing data analysis to determine power laws. The resulting relationships are $\overline{AUROC}=0.479D^{0.0568}$ and $\overline{AUPRC}=0.0157x^{0.239}$, where $D$ represents the number of paired pretraining CT imaging and report data.
Counterfactual Analysis
Methods: Counterfactual analysis (Figure 3(e)) sought to evaluate whether a model is using expected features during phenotype classification. Models can instead learn shortcuts from spurious correlations in the training data [84, 30, 20], which are easier to learn compared to more subtle physiological features that the model should use. We employ the Latent Shift approach [19] for counterfactual generation. This method uses a latent variable model, which represents the input image in a low dimensional space and reconstructs the image in its output. The gradient of the Merlin prediction is passed back to the low-dimensional space, where the latent representation is modified such that Merlin’s prediction decreases. This generates a modified image at the output of the latent variable model, which can be observed. The latent variable model regularizes the change in pixels so that the image remains close to the data manifold. The result of the Merlin counterfactual generation process is a new CT volume, which is evaluated by observing how the classifier prediction changes and how the image features driving the classification output are exaggerated or curtailed. We leverage this method to qualitatively verify that the features used during classification are consistent with what we expect the model to have learned. For this evaluation, we identify pleural effusion and splenomegaly as phenotypes with features that are well understood and can be observed in 2D slices. We select the presented volumes such that the phenotype prediction is above the decision boundary determined by AUROC curve analysis, to ensure that we are explaining a positive prediction.
Results: In Figure 11e, we apply the latent shift counterfactual method [19] to qualitatively investigate the image features that Merlin uses to perform image classification. We examine an instance of “pleural effusion” (left) where the effusion is localized to the left lung and is reduced in the counterfactual, indicating that Merlin is leveraging the features we expect. We also examine “splenomegaly” classification (right). We observe that the size of the spleen is reduced in the counterfactual relative to the original image, adding credence to the validity of imaging features that Merlin uses for image classification.
EHR Phenotypes Task
Aside from the results provided in Section 2.7, 2D and 2D-to-3D VLMs lifted offered benefits in data scarce settings with 10% training data.
2D VLM baselines: For the EHR phenotype classification task, the performance between finetuned and non-finetuned OpenCLIP and BiomedCLIP models remained similar, with a slight 2.6% increase in AUROC and 4.9% increase in AUPRC observed for all phenotypes when finetuning was applied given 10% of the training data (Figure 11b). In contrast, ResnetCLIP achieved a more significant 11.0% AUROC and 13.0% AUPRC improvement compared to finetuned OpenCLIP and BiomedCLIP at 10% training data setting (Figure 11b). These results suggest that finetuning provides a performance advantage in few-shot settings, specifically for 2D VLMs. This highlights that while finetuning 2D VLMs can enhance performance in limited data settings, their overall capacity to capture the volumetric data representations necessary for both tasks is limited.
2D-to-3D lifted VLM baselines: Larger performance gains were observed on the EHR phenotypes classification task, with a 6.8% increase in AUROC and a 15.0% increase in AUPRC for all phenotypes in the 10% training data setting (Figure 11b). These results suggest that, similar to 2D VLMs, 2D-to-3D lifted VLMs offer advantages in few-shot settings.
Figure 11: Phenotype classification extended. (a) Counterfactual analyses of pleural effusion classification (left; image from TCIA [17]) and splenomegaly classification (image from our internal test set). We annotate the zoomed in images by outlining the pathologies. The red lines border pathologies in the original images. The blue lines border pathologies in the counterfactual images. Counterfactual outlines are drawn over the original images with dotted lines and the original image outlines are also drawn over the counterfactual images with dotted lines. This allows comparing the size and shape of the pathologies between the original images and the counterfactuals, indicating that Merlin is indeed using appropriate features for image classification. (b) Average AUROC (second row) and AUPRC (third row) at 10% and 100% pretraining on the EHR phenotype classification task. Merlin pretraining results in consistently improved performance between the few-shot and fully-supervised data regimes compared to other baselines.
B.3 Cross-modal Retrieval
TaskTask MethodMethod Recall@1 Recall@8